Let X be a metric space ,then show that

(a) A subset F of X is closed iff Fc is open .

(b) A subset E of X is open iff Ec is closed .


1
Expert's answer
2022-02-24T08:24:51-0500


(a) By the definition, "F" is a closed subset of "X", iff any accumulation point of "F" belongs to "F". If "a\notin F" then let "r=\inf\{d_X(a,x):x\in F\}". Since "d_X(a,x)\geq0" for all "x", it must be "r\geq0".


Suppose "r=0". Then "\inf\{d_X(a,x):x\in F\}=0" and thus, "\forall n\in\mathbb{N}" there exists "x_n\in F" such that "d_X(a,x_n)<1\/n". Therefore, "\lim\limits_{n\to+\infty}x_n=a" and either "a\in F" or "a" is an accumulation point of "F". Since "F" is closed, in both cases "a\in F", contrary to the condition that "a\notin F".


Therefore, "r=\inf\{d_X(a,x):x\in F\}>0". In particular, this means that for all "x\in F" "d_X(a,x)\geq r". Thus, the open ball "B_r(a)" does not contain any point of "F". Therefore, "B_r(a)\subset F^c", and we proved that "F^c" contains any its point "a\in F^c" with some open ball of non-zero radius. This means that "F^c" is an open subset of "X".


Conversely, let it be given that "F^c" is an open subset of "X". We will prove that "F" is a closed subset of "X", that is, any accumulation point of "F" belongs to "F".

Let's assume the contrary: there exists a point "a" which is an accumulation point of "F" but "a\notin F". Since "a\in F^c" and "F^c" is an open subset of "X", an open ball "B_r(a)" lies inside "F^c" for some "r>0". That is, "d_X(a,x)\geq r" for all "x\in F". But since "a" is an accumulation point of "F", there exists a sequence "x_n\in F" ("n\in\mathbb{N}" ) that converges to "a". Thus, "\lim\limits_{n\to+\infty}d_X(a,x_n)=0", which contradicts to the inequality "d_X(a,x_n)\geq r". Therefore, the assumption is not valid. This implies that "F" is a closed subset of "X".


(b) The assertion "A subset E of X is open iff Ec is closed" is equivalent to ""E^c" is closed iff "E" is open." Since "E=(E^c)^c", this is a consequence of the assertion of part (a).